OpenStudy (anonymous):

Find surface area of open rectangular box 7cm length, 4cm width, 4cm high, please show work

OpenStudy (anonymous):

7*4+2*4*7+2*4*4 =28+56+32 =116cm^2

OpenStudy (anonymous):

Open box so no lid